Solve: -2x - 13 = 8x + 7
Found 2 solutions by stanbon, gwendolyn:
Answer by stanbon(75887)

Solve:
-2x - 13 = 8x + 7
--------------
10x = -20
x = -2
=============
Cheers,
Stan H.
=============

Answer by gwendolyn(128)

The first step to solving this equation is combining the like terms.
If we add 2x to both sides of the equation, we see that:
-2x-13+2x=8x+7+2x
-13=10x+7
Then, we can subtract 7 from both sides:
-13-7=10x+7-7
-20=10x
Finally, we can divide both sides by 10 to isolate the variable:
-20/10=10x/10
-2=x
So, x is equal to -2.